OpenStudy (anonymous):

Write each equation in standard form using integers. y=3x-11 y=-2x+5 y=2/3x-25/3

OpenStudy (anonymous):

Ah.. fine. y=3x-11 y-3x=3x-3x-11 -3x+y=-11 -1(-3x+y)=-1(-11) 3x-y=11 y=-2x+5 y+2x=-2x+2x+5 2x+y=5 y=2/3x-25/3 y-2/3x=2/3x-2/3x-25/3 -2/3x+y=-25/3 LCD = 3 -3(-2/3x+y)=-3(-25/3) 2x-3y=25

OpenStudy (anonymous):

Wouldn't the last one be 2x-3y=-25 because a negative times a positive is a negative?